Reports 1-1 of 1 Clear search Modify search
Detector Operation (Operations Report)
menzione - 23:03 Wednesday 23 January 2019 (44500) Print this report
Operator report - afternoon shift

Upon arrival, I found ITF not locked and the activity on SDB2 in progress.
The problem on SDB2 has been fixed at 16:11 UTC. Then the evacuation of the SDB2 minitower started. Concluded at 18:00 UTC.
The rest of the shift was dedicated to the recovery of ITF. It has been done With the help of Chiummo for Injection system, Bersanetti, Gouaty, and Allocca.
ITF very unstable, Inj OnCall intervention was needed but problem still present.

The planned activity of OptChar / flat noise investigation has been postponed.

VPM
MdVim crashed several times. restarted properly
USGS crashed several times. restarted properly
16:40 UTC PyHvac restarted
17:03 UTC TFMoni crashed / restarted

INF - ACS
at 15:41utc I set the detection lab UTA in "portata ridotta" (SDB2 activity concluded)
 

Search
×

Warning

×